OpenStudy (anonymous):

Solve the equation. x - 3x + 7 = 7

OpenStudy (anonymous):

x=0

OpenStudy (gw2011):

x-3x+7=7 x-3x=0 -2x=0 -x=0 x=0 The answer is x=0

OpenStudy (anonymous):

x-3x + 7 = 7 -2x + 7 = 7 -2x = 0 x = 0